Find the values of x for which the function f(x) = x^3-3x ,is increasing.
1
Expert's answer
2019-03-19T09:59:18-0400

The function is increasing if its derivative is positive.

"f'(x)=(x^3-3x)'=3x^2-3"

"3x^2-3>0"

"3x^2>3"

"x^2>1"

"|x|>1"

"x \in (- \infty ;-1)∪(1;+ \infty)"


Answer: "x \in (- \infty ;-1)∪(1;+ \infty)"